Identify the region or are under tge normal cyrve corresponds to the right of z is equal to 1.36


1
Expert's answer
2021-07-16T02:24:22-0400

"P(Z>1.36) = 1 -P(Z<1.36) \\
= 1 -0.9130 \\
= 0.087"